OpenStudy (anonymous):

What is the equation of the parabola with a focus at (-3, 0) and a directrix x = 3.

OpenStudy (anonymous):

i got y=1/12 x^2

OpenStudy (anonymous):

am i right???

OpenStudy (anonymous):

since the diretrix is at x=3 it means parabola is horizontal and opening to the left.

OpenStudy (anonymous):

your answer is vertical parabola opening up

OpenStudy (anonymous):

@myko thanks! what would it be so i can check it.

OpenStudy (anonymous):

for horizontal parabola just exchenge x and y and use this: equation for the parabola with the focus (0, p) and directex y = -p is y = 1/4p x2 which is equation of parabola with vertex at (0,0) so later use translation to move it to (-3,0)

OpenStudy (anonymous):

y=1/12 x^2 @myko

OpenStudy (anonymous):

\[x=ay ^{2}\] x=3=-1/(4a) 12a=-1 a=-1/12 so: \[x=-1/12y ^{2}\]